# An alternative theory of magnetic flux tubes in strong fields via axion origin photons

### v2

@@ I Introduction
The most interesting thing is that, in addition to the sunspots after the rise of the magnetic tube with $B\sim 10^{7}~G$ at all latitudes (see Fig. 9a and Fig. 10a) and the sunspots after the rise of the magnetic tube with $B\leqslant 2\cdot 10^{5}~G$ only at low latitudes (see Fig. 9b and Fig. 10c), which are the source of the double maxima of the 11-year solar cycles (see the full experimental evidence in Section II.2.2), we are interested in the problem or, more precisely, the physics of the tilt angle of Joy’s law (V. G. Ivanov, 2012; M. Dasi-Espuig et al., 2010) and how it is related to the magnetic $\nabla\rho$-pumping?
First, about the Coriolis force. According to Käpylä, P. J. et al. (2014), the rotational effect on the flow can be measured with the local Coriolis number $\hat{\Omega}\equiv Co=2\tau\Omega$, especially if $\tau$ is estimated on the basis of the theory of mixing length, which predicts values of $\hat{\Omega}$ reaching more than 10 in the deep layers (see e.g. M. Ossendrijver (2003); A. Brandenburg & K. Subramanian (2005); P. Käpylä (2011); P. Käpylä et al. (2011)). However, on the other hand, according to Käpylä, P. J. et al. (2014), the question arises of whether or not there are differential rotation bistabilities in which antisolar (stellar-like) and/or solar-like rotation depends on the initial state or launch history, or, in addition, the question of whether there are such deep layers of solar and stellar convection zones or not is still open.
In contrast to the work of Käpylä, P. J. et al. (2014), already in 2021, astrophysicist Krivodubskii (V. N. Krivodubskij, 2005, 2021) recalled (see page 72 in (V. N. Krivodubskij, 2005)) that exactly “…Modern helioseismologocal inversions (J. Schou et al., 1998; R. Howe et al., 2000; H. M. Antia, 2002; S. Basu & H. M. Antia, 2003; S. Basu & H. Antia, 2008) show that the differential rotation rate, $\Omega(r,\theta)$, in the SCZ can be divided into domains of fast (equatorial) and slow (polar) rotation. And, what is more important, the radial angular-velocity gradient $\partial\Omega/\partial r$ has opposite signs in these domains. The angular velocity increases along the radius, $\partial\Omega/\partial r>0$, at the near-equator region, and, in contrast, it decreases along the radius, $\partial\Omega/\partial r<0$, at the polar regions.” The modernity of this considered helio-seismological inversion remains to this day!
At the same time, we will show a little later that this rotating magnetic $\nabla\rho$-pumping due to the second part (slow) secondary reconnection of the $\Omega$-loop (see $(V_{rec})_{2}$ in Figs. 10b2,c) (through the primary part (fast) secondary reconnection of the $\Omega$-loop (see $(V_{rec})_{1}$ in Fig. 10b1, also in Fig.5a-d, Fig. 8b,d, and also Fig. 8c)) both in the polar region of slow rotation, where the angular velocity increases inward of the Sun ($\partial\Omega/\partial r<0$), in the equatorial region of fast rotation, where the angular velocity mainly decreases with depth ($\partial\Omega/\partial r>0$). Moreover, the direction of the magnetic $\nabla\rho$-pumping (up or down, see Fig 10b2,c) is susceptible to the sign of the factor in Eqs. (52)-(53), which depends on the polar angle (colatitude) and the behavior of the Coriolis number functions in the convection zone.
Below we show that the estimate of the Coriolis number for solar convection in the deep layer should be $\hat{\Omega}\approx 20$. Due to Eqs. (47)-(49) and the function $\phi_{n}$ for the quantity $\hat{\Omega}\approx 20$, the following values are adopted:
$\phi_{1}\cong 0.0171,~~~\phi_{2}\cong 0.0098,$ (56)
at which the radial velocity $v_{dens}^{light}$ in Eq. (43) of toroidal field transport changes the sign at the latitude $\theta^{*}=\arccos\sqrt{\varphi_{2}/\varphi_{1}}\cong 41^{\circ}$, being negative (downward) for $\theta>\theta^{*}$ and positive (upward) for $\theta<\theta^{*}$. Using (45), we find that the value of the radial velocity in the convective zone $v_{dens}^{light}$ (see Eqs. (43)-(45)) near low latitudes (e.g. $\theta^{*}=\arccos(0.985)\cong 10^{\circ}$; see also Fig. 9b) almost completely coincides with the value of the speed (46), which was previously calculated using the toroidal field $B\approx 2\cdot 10^{5}~G$ at $0.8R_{Sun}$ (see Eq. (9) and Fig. 10b,c) in the $\Re_{dens}(\beta,\varphi)$ (see Eq. (3.13) in (L. L. Kichatinov & G. Ruediger, 1992)).
@@
From here, this raises an unexpected question: In what way, in addition to the visible magnetic tube at strong fields at all latitudes (see Fig. 9a and Fig. 10a), the “invisible” magnetic tube at high latitudes (see Fig. 10b2) and the “visible” magnetic tube at low latitudes (see Fig. 10c), depending on the time $(\tau_{rec})_{2}\sim 2.5~years$ of reconnection of the horizontal field in the tachocline (which is equal to the sum of $(\tau_{n})_{max}\approx 2~years$ and the lifetime of sunspots ($\leqslant 0.5~years$) from several days to several months (see S. D’Silva & A. R. Choudhuri (1993); Y. Fan et al. (1993, 1994)) surprisingly, practically does not depend on the direction in the radial plane down or up?
@@
The answer, surprisingly, is not very simple. We know that the vertical compression (see Fig. 9a and Fig. 10b1) of the descending nearly empty anchored magnetic tube in strong fields of $\sim 10^{7}~G$ (and via convective heating $(dQ/dt)_{2}$ is associated with a horizontal “stretching” or expansion of this magnetic tube near the tachocline, where the counter-movement of magnetic fields leads to the emergence of a collisionless interaction with fast (opposite) velocities $(V_{rec})_{1}$ (and times $(\tau_{rec})_{1}$ (<<10 days)) of the first part of the fast secondary (turbulent) reconnection (see Fig. 10b1). When the vertical compression leads to horizontal stronger “stretching” of the magnetic tube in weak fields of $\geqslant 2\cdot 10^{5}~G$ (and through radiative heating $(dQ/dt)_{1}$, then the counter motion of magnetic fields (see $(V_{rec})_{1}$ for Fig. 10b1) leads to the emergence of the second part of the secondary (turbulent) reconnection with a very slow Sweet-Parker speed $(V_{rec})_{2}$ (see Fig. 10b2; also A. Lazarian & E. T. Vishniac (1999); A. Lazarian et al. (2004); G. Kowal et al. (2009); A. Lazarian et al. (2020); E. N. Parker (1957); P. A. Sweet (1958)). This means that the emerging weak fields of $\geqslant 2\cdot 10^{5}~G$ at distances from $0.80R_{Sun}$ to near $0.85R_{Sun}$ are a remarkable source of rotating magnetic $\nabla\rho$-pumping with the rise time of the magnetic tube $(\tau_{n})_{max}\sim 2.0~years$ from the tachocline to the solar surface (see Fig. 10c), and the lifetime of sunspots ($\leqslant 0.5~years$), and consequently, the time of the magnetic tube evaporation $(\tau_{rec})_{2}\leqslant 2.5~years$ and their disappearance on the solar surface.
@@
Since the study of the tendency of the tilt angle of Joy’s law is very important for understanding the evolution of the solar magnetic field, then unlike the heavy calculations of theoretical estimates of the averaged angle tilt of sunspots with increasing latitude, we can summarize our results (see Eqs. (5)-(9) in S. D’Silva & A. R. Choudhuri (1993)), which can simultaneously be expressed in the form of simple and understandable physics of Joy’s law. It is known that, according to S. D’Silva & A. R. Choudhuri (1993) and Y. Fan et al. (1993), the Coriolis force is proportional to the magnetic buoyancy velocity $v_{B}$, which can be estimated taking into account the balance between the buoyancy force (the term on the left) and the drag force (left term) and the drag force (right term):
@@
Taking into account Eqs. (18) and (19) by Y. Fan et al. (1994) and using the theoretical values of the tilt angle $\alpha$, the latitude angle $\theta$ and the radial parameter $\xi=r/R_{Sun}\sim 0.80$, one can easily obtain a simplified equation in the following form:
@@
In this regard, we are interested in an intriguing question: Where and how do sunspots appear, in which, according to M. N. Gnevyshev (1963), the maxima of the double peak in the sunspot cycle appear? And how does the first maximum – the maximum number of spots – physically differ from the second – the maximum of large spots?
@@
Below we will show that, firstly, the first maximum differs from the second maximum by being shifted in time by 2 years (see text after Eq. (55)) later than the main (first!) maximum of the solar activity cycle. Secondly, according to equation (57), when the flux tube has a large radius and a low magnetic field, then this magnetic tube, surprisingly, is the source of the formation of the tilt angle (via the Coriolis force, and not necessarily on the surface of the photosphere (see Fig. 10c)) and a large area of the spot on the surface of the Sun. Conversely, when the flux tube has a small radius and a high magnetic field, then the magnetic tube, on the one hand, is not connected to the source of the formation of the tilt angle, and on the other hand, surprisingly, has a small spot area on the surface of the Sun.
II.2.2 New physics of observed double maxima of sunspot cycles and solutions of time interval or Gnevyshev gap
@@
The phenomenon of a double peak in the sunspot cycle was discovered by M. N. Gnevyshev (1963) during a study of coronal emission and areas of active formations in the 11-year cycle 19 (M. N. Gnevyshev, 1963). First, Gnevyshev analyzed the evolution of the annual average intensity of the green coronal spectral line $\lambda 530.3$ (which serves as a direct indicator of magnetic activity and coronal heating) in five-degree heliolatitude intervals during the cycle 19 and discovered two maxima in the intensity of the corona. According to data in (A. Antalova & M. N. Gnevyshev, 1965; M. N. Gnevyshev, 1967, 1977) the first maximum of the area of sunspot groups coincides with the main maximum of the 11-year cycle for Wolf numbers, while the second maximum is associated with an increase in the number of large sunspots (the so-called maximum of sunspot intensities) (Y. I. Vitinskij et al., 1986; M. Kopecký & G. V. Kuklin, 1969). In other words, the first maximum is the maximum in the number of sunspots, and the second is the maximum of large sunspots (M. Kopecký & G. V. Kuklin, 1969). The more large spots in the cycle, the more distinct the two peaks. These features were later confirmed and for the cycle 20 (see Fig. 12), including for the solar northern and southern hemispheres separately (M. N. Gnevyshev, 1963).
@@
In contrast to the “new solar dynamo era” (see e.g. V. N. Krivodubskij (2005); Y. Fan (2021); Z. Zhang & J. Jiang (2022); Z. Zhang et al. (2023); P. Charbonneau & D. Sokoloff (2023); G. M. Vasil et al. (2024)), we used our alternative theory of magnetic flux tubes at strong fields via photons of axion origin. The essence of double maximum 11-year sunspot cycles, the source of which is the generation of a magnetic field near the bottom of the convective zone and the subsequent rise of the field from the deep layers to the surface in the photosphere, appear only at strong fields (see Fig. 9a), which generate a very “fast” (on the order of several days; see Eq. (31) $\sim 1-2~days$) first maximum (see Fig. 9a), and, as a consequence, the obligatory addition of a section at weak fields (see Fig. 10b2,c), at which it is the first maximum that is transformed into a very “slow” (on the order of two years! See $(\tau_{n})_{max}\sim 2.0~years$ second maximum (see Fig. 10b2,c)!
In other words, the very “slow” second maximum depends only on two interrelated causes: the very slow secondary reconnection (see Fig. 10b2,c) of the magnetic flux tube, which depends both at a strong magnetic field of $2\cdot 10^{5}~G$, having two sections between $0.70R_{Sun}$ and near $0.80R_{Sun}$ and between $0.85R_{Sun}$ and the photosphere (see Fig. 10b2,c), and, surprisingly, at a weak magnetic field of $2\cdot 10^{5}~G$, where the so-called effects of turbulent diamagnetic transport and rotating $\nabla\rho$-pumping appear for the first time, which provide both the process of blocking magnetic buoyancy at high latitudes (see Fig. 10b2), and at near-equatorial latitudes predetermines the process of dominance of the Coriolis force with rotating $\nabla\rho$-pumping with $\geqslant 2\cdot 10^{5}~G$ at a distance $2.3\cdot 10^{4}~km$ between $0.80R_{Sun}$ and near $0.85R_{Sun}$ (see Fig. 10c), which ultimately generates a “shift” of the arcuate ascending loop in the form of the tilt angle of Joy’s law on the solar surface (see Fig. 10c).
In this case, the key role in the proposed mechanism of double maxima is played by two “waves” of toroidal fields, in which the first wave, without the participation of secondary reconnection, first appears at all latitudes at the lower base of the CZ, and then, as a consequence, instead of the first wave without the participation of secondary reconnection, a second wave appears, which is the same primary wave, but with the participation of secondary reconnection, only at the latitudes of the equatorial surface of the Sun.
@@
$(S_{n})_{max}^{primary}/(S_{n})_{max}^{secondary}\approx\begin{cases}2&at\sim 300~sunspots\\ 1.5&at\sim 150\div 200~sunspots\\ 1&at\sim 100~sunspots\end{cases}$ (59)
and at the same time have almost the same difference between the secondary reconnection time $(\tau_{rec})_{2}\leqslant 2.5~years$ of the horizontal field in the tachocline (which is equal to the sum of $(\tau_{n})_{max}\approx 2$ and the lifetime of sunspots ($\leqslant 0.5~years$) from several days to several months (see S. K. Solanki (2003); K. Petrovay & L. van Driel-Gesztelyi (1997)), which shift by $\sim 2.0~years$,
@@
This raises an unexpected but valid question: How does the unique source of the butterfly diagram is only in the form of 11-year fundamental sunspots or is only in the form of alternating 22-year fundamental solar activity cycles?
Here we show that, as expected, the source of the butterfly diagram is found only in the form of 11-year fundamental sunspots (see M. N. Gnevyshev (1967)). Below we show our solution with the so-called latitude-time interval (or gap) between the nearest butterflies of the remarkable solar physicist M. N. Gnevyshev (1967).
@@
  • First, these are 11-year variations of sunspots (see V. D. Rusov et al. (2021)). A unique result of our model is the fact that the 11-year periods, velocities and modulations of the S102 star (see Fig. 6b in V. D. Rusov et al. (2021)) near the black hole are a significant indicator of the density modulation of the ADM halo in the fundamental plane of the Galactic center, which closely correlates with the density modulation of baryonic matter near the SMBH. If the ADM halo modulations in the GC black hole lead to ADM density modulations on the solar surface (via vertical density waves between the disk and the black hole to the solar neighborhood), then there is an “experimental” anticorrelation identity between such indicators as the 11-year ADM density modulation in the solar interior and the modulations of solar axions (or photons of axion origin), and equivalently between the modulations of solar axions (or photons of axion origin) and the sunspot cycles!
  • Secondly, the most surprising fact is that the magnetic flux tube model at strong fields of the order of $10^{7}~G$ is a consequence of the thermomagnetic EN effect in the tachocline (see Fig. C.1a in V. D. Rusov et al. (2021)), which is in our case a remarkable source of the process with a continuous transformation of the toroidal field into a poloidal one (transformation $T\rightarrow P$), but not vice versa ($P\rightarrow T$) (see Fig. 1b and Eq. (2); also Fig. C.1a in V. D. Rusov et al. (2021)).
@@
  • Thirdly, we know why we never see sunspots at the solar poles. The answer is quite simple. We know that both the tilt angle between the geometric and magnetic meridians, which are tilted by 7.25 degrees relative to the ecliptic, and the magnetic field of the northern (or southern) part of the meridional circulation, which radially corresponds to a length of about $\sim 22^{\circ}$, add up to about $\sim 30^{\circ}$. It follows clearly from this that the magnetic field of the meridional circulation part (see e.g. Fig.1b), will always slow down the magnetic field of the flux tube, which will almost never rise to the surface of the photosphere. This means that the magnetic fields of the flux tubes will only appear below about $90^{\circ}-30^{\circ}\approx 60^{\circ}$ at high latitudes. And that is precisely why no one observes sunspots at the poles of the Sun above $60^{\circ}$!
@@
  • And fourthly, we can finally observe (see Fig. 6a of V. D. Rusov et al. (2021) for Japanese x-ray telescope Yohkoh (1991–2001)) and obtain the physics of axion origin photons only in those sunspots which are practically independent of the differential rotation of the Sun in the volume butterfly diagram. Clearly, in other sunspots which depend exactly on secondary reconnection of magnetic flux tubes and, as a consequence, on the tilt angle of Joy’s law on the surface of the Sun (see Fig. 10c), there can never be occurrences of axion origin photons in sunspots due to the differential rotation of the Sun!!!
From here we now return to the physics of the Gnevyshev gap. The answer is quite simple. The Gnevyshev intervals in the butterfly diagrams are related to the 11-year modulations of the ADM density, trapped on the Sun (via vertical density waves from the disk to the solar neighborhood) and, as a result, to anticorrelated indicators of variations in the sunspot cycles or photons of axion origin. When the ADM density in the solar interior is low (or the sunspot count is high), then the Gnevyshev intervals will be short in time. Conversely, when the ADM density in the solar interior is high (or the sunspot count is low), then the Gnevyshev intervals will be longer in time. As no surprisingly, it is the presence of ADM density modulations that shows that the appearance of “our indicated” Gnevychev intervals is the result of precisely 11-year variations of butterfly cycles with sunspots!
The most interesting thing is that it is the 11-year variations of sunspot cycles or photons of axion origin (see Figs. 5a and 6b in V. D. Rusov et al. (2021)) that are anticorrelated not only with the 11-year modulations of ADM density on the Sun, but also with the 11-year modulations of ADM density in the solar neighborhood, and, in particular, on Earth (see e.g. Fig. 13).

### v3

@@ II Physics of the theory of practically empty flux tubes in strong fields of the order $\sim 10^{7}~G$
The most interesting thing is that, in addition to the sunspots after the rise of the magnetic tube with $B\sim 10^{7}~G$ at all latitudes (see Fig. 7a and Fig. 8a) and the sunspots after the rise of the magnetic tube with $B\leqslant 2\cdot 10^{5}~G$ only at low latitudes (see Fig. 7b and Fig. 8c), which are the source of the double maxima of the 11-year solar cycles (see the full experimental evidence in Section II.2.2), we are interested in the problem or, more precisely, the physics of the tilt angle of Joy’s law [111, 112] and how it is related to the magnetic $\nabla\rho$-pumping?
@@
From here, this raises an unexpected question: In what way, in addition to the visible magnetic tube at strong fields at all latitudes (see Fig. 7a and Fig. 8a), the “invisible” magnetic tube at high latitudes (see Fig. 8b2) and the “visible” magnetic tube at low latitudes (see Fig. 8c), depending on the time $(\tau_{rec})_{2}\sim 2.5~years$ of reconnection of the horizontal field in the tachocline (which is equal to the sum of $(\tau_{n})_{max}\approx 2~years$ and the lifetime of sunspots ($\leqslant 0.5~years$) from several days to several months (see [113, 69, 114]) surprisingly, practically does not depend on the direction in the radial plane down or up?
@@
The answer, surprisingly, is not very simple. We know that the vertical compression (see Fig. 7a and Fig. 8b1) of the descending nearly empty anchored magnetic tube in strong fields of $\sim 10^{7}~G$ (and via convective heating $(dQ/dt)_{2}$ is associated with a horizontal “stretching” or expansion of this magnetic tube near the tachocline, where the counter-movement of magnetic fields leads to the emergence of a collisionless interaction with fast (opposite) velocities $(V_{rec})_{1}$ (and times $(\tau_{rec})_{1}$ (<<10 days)) of the first part of the fast secondary (turbulent) reconnection (see Fig. 8b1). When the vertical compression leads to horizontal stronger “stretching” of the magnetic tube in weak fields of $\geqslant 2\cdot 10^{5}~G$ (and through radiative heating $(dQ/dt)_{1}$, then the counter motion of magnetic fields (see $(V_{rec})_{1}$ for Fig. 8b1) leads to the emergence of the second part of the secondary (turbulent) reconnection with a very slow Sweet-Parker speed $(V_{rec})_{2}$ (see Fig. 8b2; also [95, 96, 97, 98, 99, 100]). This means that the emerging weak fields of $\geqslant 2\cdot 10^{5}~G$ at distances from $0.80R_{Sun}$ to near $0.85R_{Sun}$ are a remarkable source of rotating magnetic $\nabla\rho$-pumping with the rise time of the magnetic tube $(\tau_{n})_{max}\sim 2.0~years$ from the tachocline to the solar surface (see Fig. 8c), and the lifetime of sunspots ($\leqslant 0.5~years$), and consequently, the time of the magnetic tube evaporation $(\tau_{rec})_{2}\leqslant 2.5~years$ and their disappearance on the solar surface.
@@
Since the study of the tendency of the tilt angle of Joy’s law is very important for understanding the evolution of the solar magnetic field, then unlike the heavy calculations of theoretical estimates of the averaged angle tilt of sunspots with increasing latitude, we can summarize our results (see Eqs. (5)-(9) in [113]), which can simultaneously be expressed in the form of simple and understandable physics of Joy’s law. It is known that, according to [113] and [69], the Coriolis force is proportional to the magnetic buoyancy velocity $v_{B}$, which can be estimated taking into account the balance between the buoyancy force (the term on the left) and the drag force (left term) and the drag force (right term):
@@
Taking into account Eqs. (18) and (19) by [114] and using the theoretical values of the tilt angle $\alpha$, the latitude angle $\theta$ and the radial parameter $\xi=r/R_{Sun}\sim 0.80$, one can easily obtain a simplified equation in the following form:
@@
In this regard, we are interested in an intriguing question: Where and how do sunspots appear, in which, according to [115], the maxima of the double peak in the sunspot cycle appear? And how does the first maximum – the maximum number of spots – physically differ from the second – the maximum of large spots?
@@
Below we will show that, firstly, the first maximum differs from the second maximum by being shifted in time by 2 years (see text after Eq. (53)) later than the main (first!) maximum of the solar activity cycle. Secondly, according to equation (54), when the flux tube has a large radius and a low magnetic field, then this magnetic tube, surprisingly, is the source of the formation of the tilt angle (via the Coriolis force, and not necessarily on the surface of the photosphere (see Fig. 8c)) and a large area of the spot on the surface of the Sun. Conversely, when the flux tube has a small radius and a high magnetic field, then the magnetic tube, on the one hand, is not connected to the source of the formation of the tilt angle of Joy’s law (through the Coriolis force), and on the other hand, surprisingly, has a small spot area on the surface of the Sun.
This raises a second intriguing point. One of the first solutions in our section is secondary reconnection of the MFT, which involves a turbulent diffusivity of $10^{11}~cm/s$ and a tilt angle of Joy’s law between magnetic fields ranging from $2\cdot 10^{5}~G$ (at $0.8R_{Sun}$) to $6\cdot 10^{4}~G$ (at $0.85R_{Sun}$). Moreover, due to the relationship between turbulent diffusivity and the tilt angle of Joy’s law, the direction of magnetic $\nabla\rho$-pumping (upward or downward (see Fig. 8b2,c)) is sensitive to the sign of the multiplier in equations (45)-(45), which depends on the polar angle (co-latitude) and the behavior of the Coriolis number functions in the convective zone.
It is crucial that the existence of secondary MFT reconnection generates a secondary maximum in the 11-year sunspot cycle. This is because when the primary reconnection, in which the buoyancy of a practically empty magnetic flux tube at $\sim 10^{7}~G$ rises from the tachocline to the photosphere for $\sim 2$ to $10$ days, transitions to secondary reconnection, the latter is associated with a buoyancy of practically MFT from $\sim 6\cdot 10^{4}~G$ from $0.85R_{Sun}$ (see Fig. 8c)) to the photosphere. From this we understand that since the lifetime of $\sim 2$ years for the rotating magnetic $\nabla\rho$-pumping from $2\cdot 10^{5}~G$ at $0.8R_{Sun}$ (see Fig. 8c) to $6\cdot 10^{4}~G$ is related to the lifetime of the order of $\sim 10$ days for the buoyancy of almost MFT from $\sim 6\cdot 10^{4}~G$ to the photosphere, this means that the secondary reconnection of MFT from the tachocline to the photosphere is also equal to $\sim 2$ years.
In other words, the connection between primary and secondary reconnections (or first and second maxima) is the connection of MFTs between lifetimes of the order of $\sim 2$ to $10$ days and lifetimes of the order of $\sim 2$ years, which is preserved by 11-year sunspot cycles without any dynamo.
II.2.2 Double maxima of 11-year sunspot cycles and the Gnevyshev gap
@@
The phenomenon of a double peak in the sunspot cycle was discovered by Mstislav Gnevyshev in 1963 [115] during a study of coronal emission and areas of active formations in the 11-year cycle 19 [115]. First, Gnevyshev analyzed the evolution of the annual average intensity of the green coronal spectral line $\lambda 530.3$ (which serves as a direct indicator of magnetic activity and coronal heating) in five-degree heliolatitude intervals during the cycle 19 and discovered two maxima in the intensity of the corona. According to data in [116, 117, 118], the first maximum of the area of sunspot groups coincides with the main maximum of the 11-year cycle for Wolf numbers, while the second maximum is associated with an increase in the number of large sunspots (the so-called maximum of sunspot intensities) [119, 120]. In other words, the first maximum is the maximum in the number of sunspots, and the second is the maximum of large sunspots [120]. The more large spots in the cycle, the more distinct the two peaks. These features were later confirmed and for the cycle 20 (see Fig. 9), including for the solar northern and southern hemispheres separately [115].
In contrast to the “new solar dynamo era” (see e.g. [80, 70, 121, 122, 1, 123, 2, 124]), we used our alternative theory of magnetic flux tubes at strong fields via photons of axion origin. The essence of double maximum 11-year sunspot cycles, the source of which is the generation of a magnetic field near the bottom of the convective zone and the subsequent rise of the field from the deep layers to the surface in the photosphere, manifests itself at strong fields only (see Fig. 8a, Fig. 8b2,c). This means that the two maxima are related as follows: the first maximum (the primary reconnection of the empty MFT with $10^{7}~G$ at $0.72R_{Sun}$; see Figs. 7a, 8a), which generates a very “fast” velocity to the solar surface (on the order of several days; see Eq. (31) $\sim 1-2~days$) is transformed into a secondary maximum (the secondary reconnection of the empty MFT with $10^{6}~G$ at $0.85R_{Sun}$; see Figs. 7a, 8a), which then generates a “very slow” velocity to the solar surface (on the order of two years; see Eq. (52) of $(\tau_{n})_{max}\sim 2.0~years$ for Fig. 8b2,c).
In other words, the “very slow” second maximum depends only on two interrelated causes: the very slow secondary reconnection (see Fig. 8b2,c) of the magnetic flux tube, which depends both at a strong magnetic field of $2\cdot 10^{5}~G$, having two sections between $0.7R_{Sun}$ and near $0.80R_{Sun}$, and between $0.85R_{Sun}$ and the photosphere (see Fig. 8b2,c), and, surprisingly, at a weak magnetic field of $2\cdot 10^{5}~G$, where the so-called effects of turbulent diamagnetic transport and rotating $\nabla\rho$-pumping appear for the first time, which provide both the process of blocking magnetic buoyancy at high latitudes (see Fig. 8b2), and at near-equatorial latitudes predetermines the process of dominance of the Coriolis force with rotating $\nabla\rho$-pumping with $2\cdot 10^{5}~G$ at a distance $2.3\cdot 10^{4}~km$ between $0.80R_{Sun}$ and near $0.85R_{Sun}$ (see Fig. 8c), which ultimately generates a “shift” of the arcuate ascending loop in the form of the tilt angle of Joy’s law on the solar surface (see Fig. 8c).
@@
In this case, the key role in the proposed mechanism of double maxima is played by two “waves” of toroidal fields, in which the first wave, with a strong field of $\sim 10^{7}~G$ and without the participation of secondary reconnection, first arises at all latitudes, and then, as a consequence, instead of the first wave, a second wave appears with a weak magnetic field of $2\cdot 10^{5}~G$ between $0.80R_{Sun}$ and $0.85R_{Sun}$, which is obtained from the same primary wave, but with the participation of secondary reconnection, only at the latitudes of the equatorial surface of the Sun.
@@
$(S_{n})_{max}^{primary}/(S_{n})_{max}^{secondary}\approx\begin{cases}2&at\sim 300~sunspots\\ 1.5&at\sim 150\div 200\\ 1&at\sim 100~sunspots\end{cases}$ (56)
and at the same time have almost the same difference between the secondary reconnection time $(\tau_{rec})_{2}\leqslant 2.5~years$ of the horizontal field in the tachocline (which is equal to the sum of $(\tau_{n})_{max}\approx 2$ and the lifetime of sunspots ($\leqslant 0.5~years$) from several days to several months (see [101, 83]), which shift by $\sim 2.0~years$,
@@
This raises a valid question: How is the unique source of the butterfly diagram found only in the form of 11-year variations in sunspot numbers or only in the form of alternating 22-year fundamental cycles of solar activity?
Here we show that, as expected, the source of the butterfly diagram is found only in the 11-year variations of the sunspot numbers (see [117]). Below we present an understanding of the so-called latitude-time interval (or gap) between nearby butterflies by the remarkable solar physicist Gnevyshev [117].
@@
  • First, these are 11-year variations of sunspots (see [14]). A unique result of our model is the fact that the 11-year periods, velocities and modulations of the S102 star (see Fig. 6b in [14]) near the black hole are a significant indicator of the density modulation of the ADM halo in the fundamental plane of the Galactic center, which closely correlates with the density modulation of baryonic matter near the SMBH. If the ADM halo modulations in the GC black hole lead to ADM density modulations on the solar surface (via vertical density waves between the disk and the black hole to the solar neighborhood), then there is an “experimental” anticorrelation identity between such indicators as the 11-year ADM density modulation in the solar interior and the modulations of solar axions (or photons of axion origin), and equivalently between the modulations of solar axions (or photons of axion origin) and the sunspot cycles!
@@
  • Secondly, we know why we never see sunspots at the solar poles. The answer is quite simple. We know that both the tilt angle between the geometric and magnetic meridians, which are tilted by 7.25 degrees relative to the ecliptic, and the magnetic field of the northern (or southern) part of the meridional circulation, which radially corresponds to a length of about $\sim 22^{\circ}$, add up to about $\sim 30^{\circ}$. It follows clearly from this that the magnetic field of the meridional circulation part will always slow down the magnetic field of the flux tube, which will almost never rise to the surface of the photosphere. This means that the magnetic fields of the flux tubes will only appear below about $90^{\circ}-30^{\circ}\approx 60^{\circ}$ at high latitudes. And that is precisely why no one observes sunspots at the poles of the Sun above $60^{\circ}$!
@@
  • And Thirdly, we can finally observe (see Fig. 6a of [14] for Japanese x-ray telescope Yohkoh (1991–2001)) and obtain the physics of photons of axion origin, which are associated with primary reconnection, only in those sunspots which are practically independent of the differential rotation of the Sun (when the equator rotates faster than the poles) in the butterfly diagram. Clearly, in other sunspots which depend exactly on secondary reconnection of magnetic flux tubes and, as a consequence, on the tilt angle of Joy’s law on the surface of the Sun (see Fig. 8c), there can never be occurrences of axion origin photons in sunspots due to the differential rotation of the Sun.